The patient recovery time from a particular surgical procedure is normally distributed with a mean of 5.2 days and a standard deviation of 1.6 days.

What is the 70th percentile for recovery times? (Round your answer to two decimal places.)
Answer #1

P(X < A) = P(Z < (A - mean)/standard deviation)

Mean = 5.2 days

Standard deviation = 1.6 days

Let the 70th percentile be S

P(X < S) = 0.70

P(Z < (S - 5.2)/1.6) = 0.70

Take the value of Z corresponding to 0.70 from standard normal distribution table.

(S - 5.2)/1.6 = 0.524

S = 6.04 days